# 3D-Mol: A Novel Contrastive Learning Framework for Molecular Property Prediction with 3D Information

### v2

@@ 1 Introduction
To fully understand the knowledge in unlabeled data, numerous methods based on self-supervised learning have been proposed to enhance the performance of molecular property prediction. For example, early work[5, 6, 9] employed self-supervised learning approaches for processing data represented in the Simplified Molecular Input Line Entry System (SMILES)[10]. However, SMILES is not adequate for the representation of the topological structure of molecule, making it challenging to provide reliable results. In parallel, various self-supervised learning methods based on molecular graph[15, 17, 11, 11, 12, 16, 13, 14] despite employing molecular graph data to encode the topological structure of molecule, neglects the critical three-dimensional spatial information of the molecule. Since different 3D structures may lead to dissimilar molecular properties despite having the same 2D molecular topology. As an example shown in Figure 1, Thalidomide, a sedative treatment for morning sickness in pregnant women in the 1950s, has two distinct 3D structures, R-Thalidomide and S-Thalidomide. The former has desired drug effects, while the latter has been implicated in teratogenesis. Recently, several works[18, 22, 24, 25, 22, 23, 47] utilizing molecular 3D structures have been introduced. However, limited by pretraining methods, they have not fully learned the 3D spatial information in unlabeled data. Specifically, these methods focus only on the most stable (lowest energy) 3D conformations, neglecting other existing conformations. Therefore, it is imperative to develop an approach that comprehensively acquires 3D analytical insights, encompassing both pretrain strategy and encoding technique.
To address these issues, we propose a novel framework, 3D-Mol, for molecular representation and property prediction.
@@
We employ three graphs to hierarchically represent the atom-bond, bond-angle, and dihedral information of molecule, integrating information from these hierarchies through a message-passing strategy to obtain a comprehensive molecular representation. Moreover, by using a vast amount of unlabeled data, we create a novel self-supervised method, weighted contrastive learning, to pretrain our molecular encoder alongside the geometric approach from GearNet[26]. In the proposed contrastive learning, conformations derived from the same SMILES are considered weighted positive pairs, while different ones are treated as weighted negative pairs, with weights indicating 3D conformation descriptor/fingerprint similarity. The molecular encoder is then finetuned on downstream tasks to predict molecular properties. Finally, we compare our approach with several state-of-the-art(SOTA) baselines on 7 molecular property prediction benchmarks[27], where our method achieves the best results on 5 benchmarks. In summary, our main contributions are as follows:
@@ 2.1 Molecular Representation and Encoder
The works mentioned above focus on 2D molecular representation, which might miss crucial chemical details[35] and prove insufficient for accurate molecular property prediction[li2022deep]. Recently, some studies have attempted to enhance performance by modeling 3D molecular structure. SGCN[25] applied different weights according to atomic distances during the GCN-based message passing process. SchNet[18] modeled complex atomic interactions using Gaussian radial basis functions for potential energy surface prediction to accelerate the exploration of chemical space. DimeNet[22] proposed directional message passing to fully utilize directional information within molecule. GEM[23] developed a novel geometrically enhanced molecular representation learning method and employs a specifically designed geometric-based GNN structure. However, these methods do not fully exploit the 3D structural information(like dihedral angle) of molecule.
@@ 3.1.1 Hierarchical Graph
Molecular raw data is represented by SMILES in most molecular databases. To extract spatial structure information from molecule, we use RDKit to transform the SMILES representation into 3D molecular conformations. To fully extract 3D construct information, we deconstruct molecular conformation into three hierarchical graphs, denoted as $Mol=\{G_{a-b},G_{b-a},G_{d-a}\}$. The atom-bond graph, commonly used as a 2D molecular graph, is represented as $G_{a-b}=\{V,E,P_{atom},P_{bond}\}$, where $V$ is the set of atoms and $E$ is the set of bonds. $P_{atom}\in R^{|V|*d_{atom}}$ are the attributes of atoms, and $d_{atom}$ is the number of atom attributes. $P_{bond}\in R^{|E|*d_{bond}}$ are the attributes of bonds, and $d_{bond}$ is the number of bond attributes. The bond-angle graph, is represented as $G_{b-a}=\{E,P,Ang_{\theta}\}$, where $P$ is a set of the plane that is comprised of 3 connected atoms. $Ang_{\theta}$ is the set of corresponding bond angles $\theta$. The dihedral-angle graph, is represented as $G_{d-a}=\{P,D,Ang_{\phi}\}$. The attributes of the plane are the attributes of 3 connected atoms and the corresponding bonds. $D$ represents the set of two connected planes, which connect with a bond. $Ang_{\phi}$ represents the corresponding dihedral angle $\phi$. These three graphs represent an actual molecule, and help our encoder learn 3D structure information.
@@ 3.1.2 Attribute Embedding
where $\mu^{k}_{\theta}$ and $\beta^{k}_{\theta}$ are the center and width of $\theta$, and $\mu^{k}_{\phi}$ and $\beta^{k}_{\phi}$ are the center and width of $\phi$. The centers of bond angle and dihedral angle are represented as $\mu^{k}_{\phi}$ and $\mu^{k}_{\theta}$, and the numerical values of them are $\pi$/K, where K is the number of feature dimensions.
@@ 3.1.3 Graph Embedding
where $agg^{(i)}_{d-a}$ is the aggregation function for aggregating messages from the plane neighborhood. $Comb^{(k)}_{d-a,n}$ is the update functions for updating the plane latent vector. $a^{i}_{uvw}$ is the information from the neighboring plane and the corresponding dihedral angle after being aggregated. The representation vectors of the atoms at the final iteration are integrated to gain the molecular graph representation vector $F_{mol}$ by the Readout function, which is formalized as:
@@ 3.2 Pretrain Strategy
To improve the performance of the 3D-Mol encoder, we employ contrastive learning for pretraining, categorizing conformations with identical topological structures as weighted positive pairs and contrasting ones as negatives, based on their 3D conformation descriptors and fingerprints. Inspired by GearNet[26], we also combine our pretraining method with self-supervised tasks based on physicochemical and geometric properties, and those details are in Appendix A. The overview of our pretraining method is shown in Figure 2, and the following are the details of our pretraining method.
Our objective is to facilitate the learning of the consistency and differences between molecular 3D conformations with same 2D structure, denoted as $Mconf_{i}$ and $Mconf_{j}$. To accomplish this, we employ weighted contrastive learning using a batch of molecular representations, with the loss function defined as follows:
@@
The similarity measurement between two latent vectors $F_{i}$, $F_{k}$ from a negative molecule pair ($Mconf_{i},Mconf_{k}$) is penalized by a weight coefficient $w^{fp}_{i,k}$, which computed by molecular fingerprint similarity between $Mconf_{i}$ and $Mconf_{k}$. $Sim_{FP}()$ evaluates similarity between molecular fingerprints, and $\lambda_{fp}\in[0,1]$ is the hyperparameter that determines the scale of penalty for faulty negatives. The details of the 3D conformation descriptors and fingerprint are shown in Appendix B.
4 Experiment
In this section, we conduct experiments on 7 benchmark datasets in MoleculeNet to demonstrate the effectiveness of our method for molecular property prediction. We use a large amount of unlabeled data and our pretrain strategy to pretrain our encoder, then use the downstream task to finetune the well-pretrained model and predict the molecular property. We compare it with a variety of SOTA methods and conduct several ablation studies to confirm the effectiveness of our method.
4.1 Datasets and Setup
4.1.1 Pretrain
[…]
@@ 4.1.1 Pretrain
Table 3: Ablation study. We study the performance of 3D-Mol in three scenarios: 3D-Mol, 3D-Mol without pretraining, 3D-Mol without dihedral-angle graph, then mark the best results in bold and underline the second best.
@@
$\rm 3D$-$\rm Mol_{w.o\ pretrain}$ $\rm 0.839_{0.005}$ $\rm 0.648_{0.013}$ $\rm 0.790_{0.004}$ $\rm 0.695_{0.007}$ $\rm\underline{0.807}_{0.027}$ $\rm\underline{1.667}_{0.037}$ $\rm 0.620_{0.004}$
$\rm 3D$-$\rm Mol_{w.o.dihes-angle-graph}$ $\rm\underline{0.844}_{0.004}$ $\rm\underline{0.649}_{0.006}$ $\rm\underline{0.791}_{0.005}$ $\rm\underline{0.698}_{0.006}$ $\rm 0.812_{0.015}$ $\rm 1.782_{0.007}$ $\rm\underline{0.612}_{0.005}$
4.1.2 Finetune
We use 7 molecular property prediction datasets obtained from MoleculeNet to demonstrate the effectiveness of 3D-Mol. These datasets encompass a range of biophysics, physical chemistry and physiology. The details of the datasets are in Appendix C. Following the previous work, we partition the datasets into train/validation/test sets in an 80/10/10 ratio for downstream tasks by scaffold splitting and report the mean and standard deviation by the results of 3 random seeds.
4.2 Metric
Consistent with prior studies, we adopt area under the receiver operating characteristic curve(ROC-AUC) as the evaluation metric for the classification datasets, which is a widely used metric for assessing the performance of binary classification tasks. For the regression datasets, we utilize root-mean-squared-error(RMSE) as the evaluation metric.
4.3 Result
a) Overall performance. To validate the efficacy of our method, we compare it with several baseline methods. The baseline methods are as follows: N-Gram[12] generated a graph representation by constructing node embeddings based on short walks. PretrainGNN[11] implemented several types of self-supervised learning tasks. 3D Infomax[48] maximized the mutual information between learned 3D summary vectors and the representations of a GNN. MolCLR[13] is a 2D-2D view contrastive learning model that involves atom masking, bond deletion, and subgraph removal. GraphMVP[46] used 2D-3D view contrastive learning approaches. GROVER[14] focuses on node and graph level representation and corresponding pretraining tasks for each level. GEM[23] employed predictive geometry self-supervised learning schemes that leverage 3D molecular information. Uni-Mol[47] enlarged the application scope and representation ability of molecular representation learning by using transformer. Table 1 presents compelling evidence that methods based on 3D information surpass those based on 2D in molecular modeling, as evidenced by the improved outcomes across multiple datasets. Besides, compared to current 3D-based approaches, our method achieves the best results in 5 datasets and the second-best in Tox21, highlighting its exceptional performance. Notably, our method exhibits a remarkable lead in BACE dataset. This not only affirms the value of considering spatial configurations in predictive models but also indicates our method’s potential in utilizing this information for high-precision molecular property predictions. Moreover, our method’s dominance extends across all datasets, except for some toxicity datasets. In these cases, its focus on geometric rather than substructural information, which is crucial for toxicity prediction, suggests an avenue for further refinement.

### v3

@@ 1 Introduction
To fully understand the knowledge in unlabeled data, numerous methods based on self-supervised learning have been proposed to enhance the performance of molecular property prediction. For example, early work[1, 2, 3] employed self-supervised learning approaches for processing data represented in the Simplified Molecular Input Line Entry System (SMILES)[4]. However, SMILES is not adequate for the representation of the topological structure of molecule, making it challenging to provide reliable results. In parallel, various self-supervised learning methods based on molecular graph[5, 6, 7, 8, 9, 10, 11] despite employing molecular graph data to encode the topological structure of molecule, neglects the critical three-dimensional spatial information of the molecule. Since different 3D structures may lead to dissimilar molecular properties despite having the same 2D molecular topology. As an example shown in Figure 1, Thalidomide, a sedative treatment for morning sickness in pregnant women in the 1950s, has two distinct 3D structures, R-Thalidomide and S-Thalidomide. The former has desired drug effects, while the latter has been implicated in teratogenesis. Recently, several works[12, 13, 14, 15, 13, 16, 17] utilizing molecular 3D structures have been introduced. However, limited by pretraining methods, they have not fully learned the 3D spatial information in unlabeled data. Specifically, these methods focus only on the most stable (lowest energy) 3D conformations, neglecting other existing conformations. Therefore, it is imperative to develop an approach that comprehensively acquires 3D analytical insights, encompassing both pretrain strategy and encoding technique.
To address these issues, we propose a novel framework, 3D-Mol, for molecular representation and property prediction. We employ three graphs to hierarchically represent the atom-bond, bond-angle, and dihedral information of molecule, integrating information from these hierarchies through a message-passing strategy to obtain a comprehensive molecular representation. Moreover, by using a vast amount of unlabeled data, we create a novel self-supervised method, weighted contrastive learning, to pretrain our molecular encoder alongside the geometric approach from GearNet[18].
@@
In the proposed contrastive learning, conformations derived from the same SMILES are considered weighted positive pairs, while different ones are treated as weighted negative pairs, with weights indicating 3D conformation descriptor/fingerprint similarity. The molecular encoder is then finetuned on downstream tasks to predict molecular properties. Finally, we compare our approach with several state-of-the-art(SOTA) baselines on 7 molecular property prediction benchmarks[19], where our method achieves the best results on 5 benchmarks. In summary, our main contributions are as follows:
@@ 2.1 Molecular Representation and Encoder
The works mentioned above focus on 2D molecular representation, which might miss crucial chemical details[29] and prove insufficient for accurate molecular property prediction[30]. Recently, some studies have attempted to enhance performance by modeling 3D molecular structure. Significant efforts have been made to employ 3D voxel-based representations for understanding molecular structures. Stepniewska et al. [31] used 3D convolutions to estimate the binding affinity of ligand-receptor complexes. libmolgrid [32] provided a library representing 3D molecular structures as multidimensional voxelized grids. OctSurf [33] used an octree-based representation to describe the interaction between protein pockets and ligands. In addition to voxel-based representations, several methods have been developed to embed 3D molecular information directly into GNNs. SGCN[15] applied different weights according to atomic distances during the GCN-based message passing process. SchNet[12] modeled complex atomic interactions using Gaussian radial basis functions for potential energy surface prediction to accelerate the exploration of chemical space. DimeNet[13] proposed directional message passing to fully utilize directional information within molecule. GEM[16] developed a novel geometrically enhanced molecular representation learning method and employs a specifically designed geometric-based GNN structure. However, these methods do not fully exploit the 3D structural information(like dihedral angle) of molecule.
@@ 3 Method
This section outlines the creation of 3D-Mol, a framework designed for 3D structural molecular property prediction, focusing on hierarchical graph-based molecular representation and the strategic weighting of contrastive pairs. Figure 2 provides an overview, with subsequent parts delving into specifics.
@@ 3.1.1 Hierarchical Graph
Molecular raw data is represented by SMILES in most molecular databases. To extract spatial structure information from molecule, we use RDKit[45] to transform the SMILES representation into 3D molecular conformations. To fully extract 3D construct information, we deconstruct molecular conformation into three hierarchical graphs, denoted as $Mol=\{G_{a-b},G_{b-a},G_{d-a}\}$. The atom-bond graph, commonly used as a 2D molecular graph, is represented as $G_{a-b}=\{V,E,P_{atom},P_{bond}\}$, where $V$ is the set of atoms and $E$ is the set of bonds. $P_{atom}\in R^{|V|*d_{atom}}$ are the attributes of atoms, and $d_{atom}$ is the number of atom attributes. $P_{bond}\in R^{|E|*d_{bond}}$ are the attributes of bonds, and $d_{bond}$ is the number of bond attributes. The bond-angle graph, is represented as $G_{b-a}=\{E,P,Ang_{\theta}\}$, where $P$ is a set of the plane that is comprised of 3 connected atoms. $Ang_{\theta}$ is the set of corresponding bond angles $\theta$. The dihedral-angle graph, is represented as $G_{d-a}=\{P,D,Ang_{\phi}\}$.
The attributes of the plane are the attributes of 3 connected atoms and the corresponding bonds. $D$ represents the set of two connected planes, which connect with a bond. $Ang_{\phi}$ represents the corresponding dihedral angle $\phi$. These three graphs represent an actual molecule, and help our encoder learn 3D structure information.
@@ 3.1.2 Attribute Embedding
Where $\mu^{k}_{\theta}$ and $\mu^{k}_{\phi}$ are denoted as the centers of bond angles and dihedral angles, respectively, establishing the peak of the function and centralizing the feature transformation. Similarly, the widths that determine the spread of the RBF, are represented as $\beta^{k}_{\theta}$ for bond angles and $\beta^{k}_{\phi}$ for dihedral angles. These widths dictate the spread of the function. The numerical values for these centers are set at $\pi$/K, where K is the number of feature dimensions.
@@ 3.1.3 Graph Embedding
where $agg^{(i)}_{d-a}$ is the aggregation function for aggregating messages from the plane neighborhood. $Comb^{(k)}_{d-a,n}$ is the update functions for updating the plane latent vector. $a^{i}_{uvw}$ is the information from the neighboring plane and the corresponding dihedral angle after being aggregated.
The representation vectors of the atoms at the final iteration are integrated to gain the molecular graph representation vector $F_{mol}$ by the Readout function, which is formalized as:
@@ 3.2 Pretrain Strategy
To improve the performance of the 3D-Mol encoder, we employ contrastive learning for pretraining, categorizing conformations with identical topological structures as weighted positive pairs and contrasting ones as negatives, as shown in figure 4. Inspired by GearNet[18], we also combine our pretraining method with self-supervised tasks based on physicochemical and geometric properties.
Our objective is to facilitate the learning of the consistency and differences between molecular 3D conformations. To accomplish this, we employ weighted contrastive learning using a batch of molecular representations, with the loss function defined as follows:
@@
where the two conformations with same SMILES, denoted as $Mconf_{i}$ and $Mconf_{j}$. $F_{i}$ is the latent vector extracted from $Mconf_{i}$, and $sim()$ measures the similarity between latent vectors, penalized by a weight coefficient $w^{conf}_{i,j}$, which is computed based on the 3D conformation descriptor similarity between $Mconf_{i}$ and $Mconf_{j}$. $w^{conf}_{i,j}$ represents the similarity between $Dsp_{i}$ and $Dsp_{j}$, which correspond to the 3D conformation descriptors of $Mconf_{i}$ and $Mconf_{j}$.
Figure 4: The overview of weighted positive/negative pairs. Using the molecule Oc1ccc(cc1)CC[NH3+]. a). Low weight positive pairs: Features two conformations with the same SMILES and significant structural differences (e.g., chirality, geometric angles). b). High weight positive pairs: Depicts conformations with the same SMILES but minor differences, such as slight dihedral angle variations. c). Low weight negative pair: Shows conformations from different SMILES with similar scaffolds but missing little functional group, like an (-OH) group. d). High weight negative pair: Illustrates conformations from different SMILES, differing greatly in both scaffold and functional groups.
[…]
$L_{i,j}^{l}=(f_{l}(Fn^{mk}_{n,i},Fn^{mk}_{n,j})-l_{i,j})^{2}$ (16)
where $f_{\phi}()$, $f_{\theta}()$ and $f_{l}$ are the MLPs for the local geometry task, and $L_{i,j}^{l}$, $L_{i,j,k}^{\theta}$, $L_{i,j,k,p}^{\phi}$ and $L_{i}^{FP}$ are the loss functions for each task. $CE()$ is the cross entropy loss, and $bin()$ is used to discretize the bond angle and dihedral angle. $Fn^{mk}_{n,i}$ is the latent vector of node i after masking the corresponding sampled items in each task.
$L_{i}^{FP}=BCE(f_{FP}(F^{mk}),FP_{i})$ (19)
$L_{i,j}^{dist}=(f_{dist}(Fn^{mk}_{n,i},Fn^{mk}_{n,j})-dist_{i,j})^{2}$ (20)
where $f_{FP}$ and $f_{dist}()$ are the MLPs for global geometric tasks, and $L_{i}^{FP}$ and $L_{i,j}^{dist}$ are the loss functions for each task. $BCE()$ is binary cross entropy loss. $F^{mk}$ is the latent vector of the masking molecule.
In the culmination of our pretraining stage, we consolidate the various loss functions into a unified training objective through an uncertainty weighted sum approach. The following is the final loss function:
$\begin{split}L^{final}=&L^{FP}/{\sigma_{FP}^{2}}+L^{dist}/{\sigma_{dist}^{2}}+L^{\phi}/{\sigma_{\phi}^{2}}+L^{\theta}/{\sigma_{\theta}^{2}}+L^{l}/{\sigma_{l}^{2}}+L^{conf}/{\sigma_{conf}^{2}}\\ &+log\sigma_{FP}+log\sigma_{dist}+log\sigma_{\phi}+log\sigma_{\theta}+log\sigma_{l}+log\sigma_{conf}\end{split}$ (21)
Where $L^{final}$ is the final loss we use to train the encoder in pretraining stage, and the $\sigma_{FP}$, $\sigma_{dist}$, $\sigma_{\phi}$, $\sigma_{\theta}$, $\sigma_{l}$ and $\sigma_{conf}$ are the uncertainty associated with each loss component, representing the model’s confidence in each of these loss terms. This method employs individual uncertainty terms for each loss component, allowing the model to dynamically adjust the influence of each based on its confidence in the respective predictions, which facilitates a balanced optimization across diverse molecular features, from spatial arrangements to angular orientations.
4 Experiment
In this section, we conduct experiments on 7 benchmark datasets in MoleculeNet[19] to demonstrate the effectiveness of our method for molecular property prediction. We use a large amount of unlabeled data and our pretrain strategy to pretrain our encoder, then use the downstream task to finetune the well-pretrained model and predict the molecular property. We compare it with a variety of SOTA methods and conduct several ablation studies to confirm the effectiveness of our method.
4.1 Datasets and Setup
4.1.1 Pretrain
@@
4.1.2 Finetune
We use 7 molecular property prediction datasets obtained from MoleculeNet to demonstrate the effectiveness of 3D-Mol. These datasets encompass a range of biophysics, physical chemistry and physiology. The details of the datasets are as followings:
  • BACE. The BACE dataset provides both quantitative (IC50) and qualitative (binary label) binding results for a set of inhibitors targeting human $\beta$-secretase 1 (BACE-1).
  • Tox21. The Tox21 initiative aims to advance toxicology practices in the 21st century and has created a public database containing qualitative toxicity measurements for 12 biological targets, including nuclear receptors and stress response pathways.
  • Toxcast. ToxCast, an initiative related to Tox21, offers a comprehensive collection of toxicology data obtained through in vitro high-throughput screening. It includes information from over 600 experiments and covers a large library of compounds.
  • SIDER. The SIDER database is a compilation of marketed drugs and their associated adverse drug reactions (ADRs), categorized into 27 system organ classes.
  • ESOL. The ESOL dataset is a smaller collection of water solubility data, specifically providing information on the log solubility in mols per liter for common organic small molecules.
  • FreeSolv. The FreeSolv database offers experimental and calculated hydration-free energy values for small molecules dissolved in water.
  • Lipo. Lipophilicity is a crucial characteristic of drug molecules that affects their membrane permeability and solubility. The Lipo dataset contains experimental data on the octanol/water distribution coefficient (logD at pH 7.4).
Extending from previous studies, we partition our datasets into training, validation, and test sets in an 80/10/10 ratio using scaffold splitting. This method groups molecules by their core structures, ensuring that each set features unique chemical scaffolds, in contrast to random splitting which allocates data indiscriminately of molecular similarity. This approach rigorously tests the model on novel chemical entities, offering a more stringent evaluation of its generalization capabilities. We report the mean and standard deviation by the results of 3 random seeds. The details of the finetuning settings and are in Appendix C.
4.2 Metric
In alignment with previous research, we employ the area under the receiver operating characteristic curve (ROC-AUC) as our evaluation metric for classification datasets. ROC-AUC is a prevalent and reliable measure for gauging the effectiveness of binary classification models. For regression datasets, we apply root-mean-squared-error (RMSE) as our assessment metric, which is a standard for evaluating the accuracy of regression models in predicting continuous variables.
4.3 Result
4.3.1 Overall performance
To validate the efficacy of our method, we compare it with several baseline methods. The baseline methods are as follows: N-Gram[8] generated a graph representation by constructing node embeddings based on short walks. PretrainGNN[7] implemented several types of self-supervised learning tasks. 3D Infomax[48] maximized the mutual information between learned 3D summary vectors and the representations of a GNN. MolCLR[10] is a 2D-2D view contrastive learning model that involves atom masking, bond deletion, and subgraph removal. GraphMVP[44] used 2D-3D view contrastive learning approaches. GROVER[11] focuses on node and graph level representation and corresponding pretraining tasks for each level.
@@ 4.3.1 Overall performance
GEM[16] employed predictive geometry self-supervised learning schemes that leverage 3D molecular information.
Table 3: Ablation study. We study the performance of 3D-Mol in four scenarios: 3D-Mol, 3D-Mol without pretraining, 3D-Mol without weight of contrastive learning, 3D-Mol without dihedral-angle graph, then mark the best results in bold and underline the second best.
@@
$\rm 3D$-$\rm Mol_{w.o\ pretrain}$ $\rm 0.839_{0.005}$ $\rm 0.648_{0.013}$ $\rm 0.790_{0.004}$ $\rm 0.695_{0.007}$ $\rm 0.807_{0.027}$ $\rm\underline{1.667}_{0.037}$ $\rm 0.620_{0.004}$
$\rm 3D$-$\rm Mol_{w.o.cl-weight}$ $\rm\underline{0.851}_{0.003}$ $\rm 0.645_{0.009}$ $\rm 0.786_{0.005}$ $\rm 0.696_{0.002}$ $\rm\underline{0.795}_{0.016}$ $\rm 1.705_{0.038}$ $\rm\underline{0.612}_{0.010}$
$\rm 3D$-$\rm Mol_{w.o.dihes-angle-graph}$ $\rm 0.844_{0.004}$ $\rm\underline{0.649}_{0.006}$ $\rm\underline{0.791}_{0.005}$ $\rm\underline{0.698}_{0.006}$ $\rm 0.812_{0.015}$ $\rm 1.782_{0.007}$ $\rm\underline{0.612}_{0.005}$
Uni-Mol[17] enlarged the application scope and representation ability of molecular representation learning by using transformer. Table 1 presents compelling evidence that methods based on 3D information surpass those based on 2D in molecular modeling, as evidenced by the improved outcomes across multiple datasets. Besides, compared to current 3D-based approaches, our method achieves the best results in 5 datasets and the second-best in Tox21, highlighting its exceptional performance. Notably, our method exhibits a remarkable lead in BACE dataset. This not only affirms the value of considering spatial configurations in predictive models but also indicates our method’s potential in utilizing this information for high-precision molecular property predictions. Moreover, our method’s dominance extends across all datasets, except for some toxicity datasets. In these cases, its focus on geometric rather than substructural information, which is crucial for toxicity prediction, suggests an avenue for further refinement.
4.3.2 Encoder performance